One school survey showed that three out of five students own a pet. Another survey showed that 6 out of 11 students onto pet. Are these results equivalent explain the reasoning
1
Expert's answer
2020-10-18T17:53:08-0400

1) In the first survey percent of those who own a pet: 3 / 5 * 100 = 60%

or "\dfrac{3}{5} * 2 = \dfrac{3*2}{5*2} =\dfrac{6}{10}".

2) And in the second: 6 / 11 * 100 ≈ 54.5%.

3) "\dfrac{6}{10}" is not equivalent to "\dfrac{6}{11}" .

Answer: No.